# Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators

### v1

@@
Directional and contra-directional coupling in Huygens’ metawaveguide microring resonators
@@
††journal: opticajournal††articletype: Research ArticleHuygens’ metawaveguides represent a transformative concept in photonic device engineering, enabling unprecedented control over light propagation. This study presents, for the first time, integrated Huygens’-based microring resonators and directional and contra-directional couplers, specifically designed for operation at the 1550 nm telecommunication wavelength. By leveraging the unique properties of resonant Huygens’ waveguides, we demonstrate efficient evanescent directional coupling with high- resonators, characterized by negative group index and near-zero dispersion, which are critical for enhancing performance in compact, high-performance add-drop filters. The research further explores the implications of these novel structures on group index and group velocity dispersion, providing insights into their potential applications in nonlinear optics and quantum information technologies. Notably, the introduction of a hybrid subwavelength grating-Huygens’ contra-directional coupler facilitates backward coupling between resonant and non-resonant metawaveguides, achieving a broad spectral rejection bandwidth. Our findings advance the integration of resonant metamaterials into scalable photonic platforms, laying the groundwork for innovative applications in optical communications, quantum photonics and sensing systems.
1 Introduction
@@
Photonic metamaterials have fundamentally transformed our capacity to manipulate both free-space and guided optical waves, enabling functionalities far beyond those achievable with conventional dielectric media. Through precise engineering of their optical parameters, metamaterials provide unprecedented control over key properties of light—such as amplitude, phase, polarization, wavelength, and group velocity [1, 2, 3, 4, 5, 6, 7]. This control is typically realized by subwavelength-scale periodic modulation of materials with contrasting refractive indices [7, 8, 9, 10, 11, 12, 13]. Moreover, by harnessing electromagnetic resonances within these periodic nanostructures, the light–matter interaction can be significantly amplified [14, 15, 16, 17].
@@
Resonant metamaterials exhibit a range of distinctive functionalities, including tailored light transmission [18], enhanced absorption [19], control over propagation direction [20, 21], beam shaping, focusing and lensing [22, 23], sensing [24], and enhancement of nonlinear optical effects [25]. These capabilities are vital for diverse applications spanning classical and quantum information processing, communication, and signal routing [26, 27, 28].
@@
Manipulation of the effective refractive index and group index is essential for optimizing mode coupling [29, 30] and managing pulse propagation and dispersion in integrated photonics [31, 32, 33]. Such control enables advanced signal processing techniques indispensable to high-performance optical networks. Dispersion-engineered metamaterials, especially those incorporating subwavelength grating (SWG) structures that modulate the effective index [34, 35, 36], have emerged as key components in fiber-optic communication systems. These structures facilitate efficient fiber-to-chip coupling [37, 38],mode conversion [39], wavelength multiplexing [37], and high-performance directional and contra-directional couplers [40, 41, 42, 43, 44, 45], as well as micro-resonators [46, 47, 48, 49] and other advanced photonic devices [12, 13].
@@
In parallel, recent studies have unveiled remarkable optical phenomena — such as backward pulse propagation [50, 51, 52, 53, 54, 55], pulse compression [56, 57, 58, 59], ultrafast signal processing and switching [60, 61, 62], and soliton and frequency-comb generation [63, 64, 65] — enabled by waveguides exhibiting negative group index and anomalous dispersion [66, 67, 68]. Negative group index, wherein the phase velocity is opposite to the energy flow, gives rise to counterintuitive phenomena such as reversed pulse propagation and enhanced control of optical information. These regimes are accessible in resonant metamaterials [69, 70].
@@
While early implementations of resonant metawaveguides based on periodic Mie-scattering nanoparticles were hindered by excessive scattering and absorption losses [7, 70, 14, 71], recent progress has led to the realization of low-loss Huygens’ metawaveguides [72, 73]. These structures exploit Huygens’ particles, which achieve backscattering suppression through constructive interference between co-aligned electric dipole (ED) and magnetic dipole (MD) resonances—satisfying the Kerker condition [74, 75, 18, 76, 77]. This condition ensures forward-directed radiation, enabling light propagation via near-field coupling between adjacent Huygens’ particles, while preventing scattering into radiation modes.
@@
2 Results and Discussions
2.1 Huygens’ nanoantennas and metawaveguides
The core components of Huygens’ metawaveguides are individual dielectric nanoantennas that simultaneously exhibit Mie resonant electric and magnetic dipole (ED and MD) scattering within a desired spectral range [72]. In Fig. 1a(i) the ED and MD resonances in a Huygens particle are shown schematically, with the respective moments and satisfying the forward Kerker condition, (where is the medium’s relative permittivity and is the speed of light in vacuum), ensuring their comparable magnitudes and spectral overlap [78, 79]. This in-phase excitation creates constructive interference of the ED and MD scattering in the forward direction and destructive interference in the backward direction, hence in zero backward scattering. Under this condition, the nanoantennas function as forward directional Huygens’ scatterers.
Through multipolar decomposition analysis of the total scattering cross-section [80] (see Fig. S1 in Supplement 1), we identified that single-crystal silicon nano-cuboid antennas with dimensions nm, nm, and nm embedded in a silica cladding act as unidirectional forward Huygens’ scatterers at the 1550 nm telecommunication wavelength. This is confirmed by the near-field mode profile shown in Fig. 1a(ii). The height nm was specifically chosen to ensure compatibility with the standard CMOS compatible silicon-on-insulator (SOI) waveguide platform. We carefully tuned the cuboid width-to-length ratio, , for TE polarization to achieve spectral overlap of the ED and MD resonances near the 1550 nm wavelength. We carried out the multipolar expansion of the individual Huygens’ nanoantennas by exciting it with a totalfield scattered-field (TFSF) source in a commercial software Ansys Lumerical FDTD, and performed the decomposition of the total scattering cross-section using the generalized Kerker condition calculation.
@@ 2.1 Huygens’ nanoantennas and metawaveguides
The calculated dispersion band of the infinitely long metawaveguide (see Fig. 1(c)) reveals two guided modes separated by the photonic band gap (PBG) around 1600 nm. The light line of SiO2 is indicated in yellow. The fundamental guided TE mode lies primarily in the long-wavelength region (below the PBG), referred to as the SWG band. Here, the chosen nanoantenna dimensions are considerably smaller than the wavelength of the propagating light. The SWG band extends up to the PBG at the edge of the Brillouin zone, where the array period equals half the effective wavelength. Below the PBG, the behavior of the SWG mode can be described by expressing the effective material refractive index, , as a weighted average of the constituent refractive indices. The second guided mode, positioned above the PBG near the Huygens’ nanoantenna resonance wavelength, operates in the Huygens’ band and exhibits a negative group index [72].
We chose the antenna periodicity, = 430 nm, to spectrally position the Huygens’ band in the 1550 nm telecommunication wavelength range and calculated the Bloch modes for a discrete set of wavevectors in normalized ranging from 0.3 to 0.5 (2). The chosen periodicity, of individual antennas with width, constitutes the duty cycle, = 0.73. Next, we calculated the field profiles of the Bloch modes ( from the middle of the – cross-section of the antenna) of each branch at = 0.425 and 1550 nm for the chosen periodicity, and launch the Bloch mode through the straight Huygens’ metawaveguide. We performed a multipolar decomposition analysis of the fields excited inside the individual antennas for the 1D metawaveguide periodic array. Multipole decomposition for the Huygens’ band revealed a good spectral overlap between ED and MD resonances, as required to fulfill the Kerker condition over a broadband (see Fig. S1 in Supplement 1). The resonant properties of the coupled individual antennas are preserved in the array while the relative magnitudes of the electric and magnetic moments are locked across the Huygens band. The antennas in the array effectively act as coupled Huygens’ scatterers over a wide spectral range of 60 nm [72].
@@
The fabricated samples were characterized using a continuous wave (CW) laser Agilent 8164B Lightwave Measurement System (LMS) with a linewidth of 0.01 nm operating in the wavelength range of 1460 nm to 1600 nm. To launch the input and collect and output light to measure the transmission, we used two spherical lensed fibers (radius of curvature: = 9 m and mode field diameter: MFD = 3 m); one was connected to the TE polarization-maintained laser source output, and another to the power meter. Every time we measured the transmission, we scanned the CW laser across its operating wavelength range to cover the transmission spectra of the ring resonators. The input and outputs of the bus waveguides were directly connected to the on-chip fiber-to-edge couplers [37] to couple the source light and collect the transmitted signal.
2.2 Directional coupling of Huygens’ metawaveguides
@@
We designed and fabricated on-chip SOI directional couplers by placing two straight Huygens’ metawaveguides in parallel, separated by a coupling gap, , to enable evanescent mode coupling, and finally measured the transmission power over the entire Huygens’ band from 1480 nm to 1560 nm wavelength for all the devices. The mode propagation in the directional coupler and the corresponding scanning electron microscopy (SEM) image of the device are shown in Fig.1(d) and Fig.1(e), respectively. The coupling length was designed as m and the duty cycle to to achieve 100 coupling efficiency [46].
@@
To evaluate the impact of the coupling gap on the propagation characteristics of the coupled Huygens’ metawaveguides, we conducted a multipolar expansion analysis of individual antennas at the central location of the coupler. Fig.1(f) presents the simulated scattering cross-section for decomposed multipolar components with a minimum coupling gap of 100 nm. While the electric dipole (ED) and magnetic dipole (MD) are excited in phase near 1576 nm, they exhibit an amplitude mismatch of nearly a factor of 2, yet remain within the same order of magnitude. The amplitude matching is improved for increased coupling gaps as shown in Fig.1(g), while the spectral overlap remains consistent (see Fig. S1 in Supplement 1). Fig. 1(h) reveals that the coupling efficiency of the straight Huygens’ directional couplers strongly depends on the gap between the adjacent waveguides. Experimental results arepresented in Fig. S2 in Supplement 1. This behavior is expected since variations in the gap affect significantly the mode index of the Huygens’ waveguide in the coupling region.
@@
Since Huygens’ scatterers are designed to suppress back-scattering (see Fig.1(a)), they effectively reduce propagation losses associated with back-reflections in periodic resonant metawaveguides. Our results confirm that back-reflection remains suppressed even in coupled Huygens’ metawaveguides, where energy exchange between grating structures could otherwise induce back-scattering [81]. Comparisons with a non-resonant SWG metawaveguide directional coupler with the same coupling length ( = 9 m), duty cycle ( = 0.72), antenna width ( = 180 nm), and periodicity ( = 250 nm) show that backward directional coupling remains lower in the coupled Huygens’ waveguides compared to the SWG waveguides, as shown in Fig.1(i).
2.3 Huygens’ ring resonators
@@
Using the FSR corresponding to the resonances shown in Fig.2(e), we extracted the metawaveguide group index, , from the relation , where is the speed of light in vacuum and is the length of the ring. Due to the inverse frequency-wavevector dispersion relation (shown in Fig.1(c)), the group index, , for the Huygens’ band is negative. The flat and descending FSR depicted in Fig.2(e) correspond to zero and anomalous group velocity dispersion (GVD), respectively. The dispersion parameter, , where , also follows this behavior. Fig.2(f) depicts the group index, , averaging -4.65 in the central zone of the Huygens’ band, decreasing to -10 near the PBG edges. In contrast, the dispersion parameter, , in the same plot demonstrates near zero dispersion in the central zone which eventually declines to -10,000 ps/nm·km as the photonic bandgap is approached. Such feature is unique for the Huygens’ metawaveguide ring resonators, as shown in the comparison with the conventional and the SWG ring resonators presented in Fig. S3 in Supplement 1.
2.4 Characteristics of Huygens’ micro-resonator add-drop filters
@@
In the ring-based filters, the observed resonances exhibit over-coupled, critically coupled, and under-coupled characteristics as the coupling gap increases. The extinction ratio of the through-port resonances gradually decreases with increasing coupling gap, as the coupling strength is reduced. For both over- and critically coupled regimes, the insertion loss difference between the add and drop ports is approximately 10 dB. In the under-coupled regime, some resonances exhibit spectral splitting [82, 81, 83], accompanied by a reduced insertion loss difference between the add and drop ports. The resonance splitting would become more pronounced for larger coupling gaps, as shown in Fig. S4 in Supplement 1.
In the racetrack filter, the resonances transition from over-coupled to critically coupled and then to under-coupled as the coupling gap increases. For the smallest coupling gap, the over-coupled resonance extinction ratio at the through port is approximately 5–10 dB. This value increases to 30–40 dB for critically coupled resonances when the gap is increased by about 50 nm. With an additional 50 nm increment, under-coupling becomes pronounced and the through-port transmission drops to 20 dB below the drop-port transmission. Unlike the ring resonators, here the resonances are split in the over-coupled regime with the spectral separation exceeding the 3-dB bandwidth. Consequently, the insertion loss at the add and drop ports becomes comparable. The asymmetric split resonances suggest back-coupling effects in the Huygens’ racetrack resonators [82, 81], which may arise from the waveguide bends or the directional coupler sections.
@@
2.5 Add-drop filtering with contra-directional couplers
@@
To conclude this section, we experimentally assessed the behaviour of the designed apodized CDC when integrated with a racetrack resonator. The device under test is shown in the SEM images of Fig. 6(a). The coupling length was varied from 20 m to 300 m in several steps, which in turn modified the free spectral range (FSR) of the resonances from 2.3 to 0.5 nm, respectively. A more detailed view of the racetrack resonances is provided in Fig. S5 in Supplement 1. The transmission spectra of the CDC-loaded racetrack resonators at the through, add and drop ports are presented in Fig. 6(b-g). The observed resonances exhibited an exceptionally low FSR, while the CDC also limited the light coupling spectrum of the racetrack to a 10-nm rejection band centered within the Huygens’ operation bandwidth. This demonstrated the ability to tailor the Huygens’ racetrack resonances with using an envelope function provided by an apodized SWG-assisted CDC.
3 Conclusion

### v2

@@
Abstract
@@
Huygens’ metawaveguides represent a transformative concept in photonic device engineering, enabling unprecedented control over light propagation. This study presents, for the first time, integrated Huygens’-based microring resonators and directional and contra-directional couplers, specifically designed for operation at the 1550 nm telecommunication wavelength. By leveraging the unique properties of resonant Huygens’ waveguides, we demonstrate efficient evanescent directional coupling with high- resonators, characterized by negative group index and near-zero dispersion, which are critical for enhancing performance in compact, high-performance add-drop filters. The research further explores the implications of these novel structures on group index and group velocity dispersion, providing insights into their potential applications in nonlinear optics and quantum information technologies. Notably, the introduction of a hybrid subwavelength grating-Huygens’ contra-directional coupler facilitates backward coupling between resonant and non-resonant metawaveguides, achieving a broad spectral rejection bandwidth. Our findings advance the integration of resonant metamaterials into scalable photonic platforms, laying the groundwork for innovative applications in optical communications, quantum photonics and sensing systems.
Keywords
Huygens’ scattering, dielectric nanoparticles, metawaveguides, integrated photonics, directional couplers, ring resonators
Introduction
@@
Photonic metamaterials have fundamentally transformed our capacity to manipulate both free-space and guided optical waves, enabling functionalities far beyond those achievable with conventional dielectric media. Through precise engineering of their optical parameters, metamaterials provide unprecedented control over key properties of light—such as amplitude, phase, polarization, wavelength, and group velocity zheludev2012metamaterials, smith2004metamaterials, shalaev2007optical, jahani2016all, yao2014plasmonic, monticone2017metamaterial, brener2019dielectric. This control is typically realized by subwavelength-scale periodic modulation of materials with contrasting refractive indices brener2019dielectric, soukoulis2011past, liu2008three, kadic20193d, chen2020flat, cheben2023recent, cheben2018subwavelength. Moreover, by harnessing electromagnetic resonances within these periodic nanostructures, the light–matter interaction can be significantly amplified kuznetsov2016optically, koshelev2020dielectric, staude2017metamaterial, bin2021ultra.
@@
Resonant metamaterials exhibit a range of distinctive functionalities, including tailored light transmission decker2015high, enhanced absorption landy2008perfect, control over propagation direction park2021all, yu2011light, beam shaping, focusing and lensing kundtz2010extreme, khorasaninejad2016metalenses, sensing chen2012metamaterials, and enhancement of nonlinear optical effects lapine2014colloquium. These capabilities are vital for diverse applications spanning classical and quantum information processing, communication, and signal routing cortes2012quantum, stav2018quantum, solntsev2021metasurfaces.
@@
Manipulation of the effective refractive index and group index is essential for optimizing mode coupling son2018high, dulkeith2006group and managing pulse propagation and dispersion in integrated photonics gauthier2005slow, boyd2009controlling, turner2006tailored. Such control enables advanced signal processing techniques indispensable to high-performance optical networks. Dispersion-engineered metamaterials, especially those incorporating subwavelength grating (SWG) structures that modulate the effective index halir2015waveguide, halir2018subwavelength, bock2010subwavelength, have emerged as key components in fiber-optic communication systems. These structures facilitate efficient fiber-to-chip coupling cheben2010refractive, papes2016fiber,mode conversion cheben2006subwavelength, wavelength multiplexing cheben2010refractive, and high-performance directional and contra-directional couplers halir2012colorless, lu2015broadband, wang2010uniform, shi2013silicon, shi2013ultra, naghdi2016silicon, as well as micro-resonators donzella2014sub, donzella2015design, eid2016fsr, naraine2023subwavelength and other advanced photonic devices cheben2023recent, cheben2018subwavelength.
@@
In parallel, recent studies have unveiled remarkable optical phenomena — such as backward pulse propagation liu2002superluminal, longhi2003superluminal, yanik2004time, stefaniuk2023nonlocality, he2022fast, zhang2021extraordinary, pulse compression peccianti2010subpicosecond, tan2010monolithic, blanco2014observation, tan2015nonlinear, ultrafast signal processing and switching chai2017ultrafast, wang2018chip, raja2021ultrafast, and soliton and frequency-comb generation del2007optical, brasch2016photonic, chang2022integrated — enabled by waveguides exhibiting negative group index and anomalous dispersion dogariu2001transparent, godey2014stability, kim2017dispersion. Negative group index, wherein the phase velocity is opposite to the energy flow, gives rise to counterintuitive phenomena such as reversed pulse propagation and enhanced control of optical information. These regimes are accessible in resonant metamaterials shcherbakov2015ultrafast, savelev2016solitary.
@@
While early implementations of resonant metawaveguides based on periodic Mie-scattering nanoparticles were hindered by excessive scattering and absorption losses brener2019dielectric, savelev2016solitary, kuznetsov2016optically, bakker2017resonant, recent progress has led to the realization of low-loss Huygens’ metawaveguides sirmaci2023all, sirmaci2025ultracompact. These structures exploit Huygens’ particles, which achieve backscattering suppression through constructive interference between co-aligned electric dipole (ED) and magnetic dipole (MD) resonances—satisfying the Kerker condition kruk2017functional, staude2013tailoring, decker2015high, terekhov2017multipolar, kerker1983electromagnetic. This condition ensures forward-directed radiation, enabling light propagation via near-field coupling between adjacent Huygens’ particles, while preventing scattering into radiation modes.
@@
Results and Discussions
Huygens’ nanoantennas and metawaveguides
The core components of Huygens’ metawaveguides are individual dielectric nanoantennas that simultaneously exhibit Mie resonant electric and magnetic dipole (ED and MD) scattering within a desired spectral range sirmaci2023all. In Fig. 1a(i) the ED and MD resonances in a Huygens particle are shown schematically, with the respective moments and satisfying the forward Kerker condition, (where is the medium’s relative permittivity and is the speed of light in vacuum), ensuring their comparable magnitudes and spectral overlap liu2018generalized, shamkhi2018transverse. This in-phase excitation creates constructive interference of the ED and MD scattering in the forward direction and destructive interference in the backward direction, hence in zero backward scattering. Under this condition, the nanoantennas function as forward directional Huygens’ scatterers.
Through multipolar decomposition analysis of the total scattering cross-section alaee2015generalized (see Fig. S1 in Supporting Information Section I), we identified that single-crystal silicon nano-cuboid antennas with dimensions nm, nm, and nm embedded in a silica cladding act as unidirectional forward Huygens’ scatterers at the 1550 nm telecommunication wavelength. This is confirmed by the near-field mode profile shown in Fig. 1a(ii). The height nm was specifically chosen to ensure compatibility with the standard CMOS compatible silicon-on-insulator (SOI) waveguide platform. We carefully tuned the cuboid width-to-length ratio, , for TE polarization to achieve spectral overlap of the ED and MD resonances near the 1550 nm wavelength. We carried out the multipolar expansion of the individual Huygens’ nanoantennas by exciting it with a totalfield scattered-field (TFSF) source in a commercial software Ansys Lumerical FDTD, and performed the decomposition of the total scattering cross-section using the generalized Kerker condition calculation.
@@
The calculated dispersion band of the infinitely long metawaveguide (see Fig. 1(c)) reveals two guided modes separated by the photonic band gap (PBG) around 1600 nm. The light line of SiO2 is indicated in yellow. The fundamental guided TE mode lies primarily in the long-wavelength region (below the PBG), referred to as the SWG band. Here, the chosen nanoantenna dimensions are considerably smaller than the wavelength of the propagating light. The SWG band extends up to the PBG at the edge of the Brillouin zone, where the array period equals half the effective wavelength. Below the PBG, the behavior of the SWG mode can be described by expressing the effective material refractive index, , as a weighted average of the constituent refractive indices. The second guided mode, positioned above the PBG near the Huygens’ nanoantenna resonance wavelength, operates in the Huygens’ band and exhibits a negative group index sirmaci2023all.
We chose the antenna periodicity, = 430 nm, to spectrally position the Huygens’ band in the 1550 nm telecommunication wavelength range and calculated the Bloch modes for a discrete set of wavevectors in normalized ranging from 0.3 to 0.5 (2). The chosen periodicity, of individual antennas with width, constitutes the duty cycle, = 0.73. Next, we calculated the field profiles of the Bloch modes ( from the middle of the – cross-section of the antenna) of each branch at = 0.425 and 1550 nm for the chosen periodicity, and launch the Bloch mode through the straight Huygens’ metawaveguide. We performed a multipolar decomposition analysis of the fields excited inside the individual antennas for the 1D metawaveguide periodic array. Multipole decomposition for the Huygens’ band revealed a good spectral overlap between ED and MD resonances, as required to fulfill the Kerker condition over a broadband (see Fig. S1 in Supporting Information Section I). The resonant properties of the coupled individual antennas are preserved in the array while the relative magnitudes of the electric and magnetic moments are locked across the Huygens band. The antennas in the array effectively act as coupled Huygens’ scatterers over a wide spectral range of 60 nm sirmaci2023all.
@@
The fabricated samples were characterized using a continuous wave (CW) laser Agilent 8164B Lightwave Measurement System (LMS) with a linewidth of 0.01 nm operating in the wavelength range of 1460 nm to 1600 nm. To launch the input and collect and output light to measure the transmission, we used two spherical lensed fibers (radius of curvature: = 9 m and mode field diameter: MFD = 3 m); one was connected to the TE polarization-maintained laser source output, and another to the power meter. Every time we measured the transmission, we scanned the CW laser across its operating wavelength range to cover the transmission spectra of the ring resonators. The input and outputs of the bus waveguides were directly connected to the on-chip fiber-to-edge couplers cheben2010refractive to couple the source light and collect the transmitted signal.
Directional coupling of Huygens’ metawaveguides
@@
We designed and fabricated on-chip SOI directional couplers by placing two straight Huygens’ metawaveguides in parallel, separated by a coupling gap, , to enable evanescent mode coupling, and finally measured the transmission power over the entire Huygens’ band from 1480 nm to 1560 nm wavelength for all the devices. The mode propagation in the directional coupler and the corresponding scanning electron microscopy (SEM) image of the device are shown in Fig.1(d) and Fig.1(e), respectively. The coupling length was designed as m and the duty cycle to to achieve 100 coupling efficiency donzella2014sub.
@@
To evaluate the impact of the coupling gap on the propagation characteristics of the coupled Huygens’ metawaveguides, we conducted a multipolar expansion analysis of individual antennas at the central location of the coupler. Fig.1(f) presents the simulated scattering cross-section for decomposed multipolar components with a minimum coupling gap of 100 nm. While the electric dipole (ED) and magnetic dipole (MD) are excited in phase near 1576 nm, they exhibit an amplitude mismatch of nearly a factor of 2, yet remain within the same order of magnitude. The amplitude matching is improved for increased coupling gaps as shown in Fig.1(g), while the spectral overlap remains consistent (see Fig. S1 in Supporting Information Section I). Fig. 1(h) reveals that the coupling efficiency of the straight Huygens’ directional couplers strongly depends on the gap between the adjacent waveguides. Experimental results are presented in Fig. S2 in Supporting Information Section II. This behavior is expected since variations in the gap affect significantly the mode index of the Huygens’ waveguide in the coupling region.
@@
Since Huygens’ scatterers are designed to suppress back-scattering (see Fig.1(a)), they effectively reduce propagation losses associated with back-reflections in periodic resonant metawaveguides. Our results confirm that back-reflection remains suppressed even in coupled Huygens’ metawaveguides, where energy exchange between grating structures could otherwise induce back-scattering li2016backscattering. Comparisons with a non-resonant SWG metawaveguide directional coupler with the same coupling length ( = 9 m), duty cycle ( = 0.72), antenna width ( = 180 nm), and periodicity ( = 250 nm) show that backward directional coupling remains lower in the coupled Huygens’ waveguides compared to the SWG waveguides, as shown in Fig.1(i).
Huygens’ ring resonators
@@
Using the FSR corresponding to the resonances shown in Fig.2(e), we extracted the metawaveguide group index, , from the relation , where is the speed of light in vacuum and is the length of the ring. Due to the inverse frequency-wavevector dispersion relation (shown in Fig.1(c)), the group index, , for the Huygens’ band is negative. The flat and descending FSR depicted in Fig.2(e) correspond to zero and anomalous group velocity dispersion (GVD), respectively. The dispersion parameter, , where , also follows this behavior. Fig.2(f) depicts the group index, , averaging -4.65 in the central zone of the Huygens’ band, decreasing to -10 near the PBG edges. In contrast, the dispersion parameter, , in the same plot demonstrates near zero dispersion in the central zone which eventually declines to -10,000 ps/nm·km as the photonic bandgap is approached. Such feature is unique for the Huygens’ metawaveguide ring resonators, as shown in the comparison with the conventional and the SWG ring resonators presented in Fig. S3 in Supporting Information Section III.
Characteristics of Huygens’ micro-resonator add-drop filters
@@
In the ring-based filters, the observed resonances exhibit over-coupled, critically coupled, and under-coupled characteristics as the coupling gap increases. The extinction ratio of the through-port resonances gradually decreases with increasing coupling gap, as the coupling strength is reduced. For both over- and critically coupled regimes, the insertion loss difference between the add and drop ports is approximately 10 dB. In the under-coupled regime, some resonances exhibit spectral splitting bogaerts2012silicon, li2016backscattering, zhang2008resonance, accompanied by a reduced insertion loss difference between the add and drop ports. The resonance splitting would become more pronounced for larger coupling gaps, as shown in Fig. S4 in Supporting Information Section IV.
In the racetrack filter, the resonances transition from over-coupled to critically coupled and then to under-coupled as the coupling gap increases. For the smallest coupling gap, the over-coupled resonance extinction ratio at the through port is approximately 5–10 dB. This value increases to 30–40 dB for critically coupled resonances when the gap is increased by about 50 nm. With an additional 50 nm increment, under-coupling becomes pronounced and the through-port transmission drops to 20 dB below the drop-port transmission. Unlike the ring resonators, here the resonances are split in the over-coupled regime with the spectral separation exceeding the 3-dB bandwidth. Consequently, the insertion loss at the add and drop ports becomes comparable. The asymmetric split resonances suggest back-coupling effects in the Huygens’ racetrack resonators bogaerts2012silicon, li2016backscattering, which may arise from the waveguide bends or the directional coupler sections.
@@
Add-drop filtering with contra-directional couplers
@@
To conclude this section, we experimentally assessed the behaviour of the designed apodized CDC when integrated with a racetrack resonator. The device under test is shown in the SEM images of Fig. 6(a). The coupling length was varied from 20 m to 300 m in several steps, which in turn modified the free spectral range (FSR) of the resonances from 2.3 to 0.5 nm, respectively. A more detailed view of the racetrack resonances is provided in Fig. S5 in Supporting Information Section V. The transmission spectra of the CDC-loaded racetrack resonators at the through, add and drop ports are presented in Fig. 6(b-g). The observed resonances exhibited an exceptionally low FSR, while the CDC also limited the light coupling spectrum of the racetrack to a 10-nm rejection band centered within the Huygens’ operation bandwidth. This demonstrated the ability to tailor the Huygens’ racetrack resonances with using an envelope function provided by an apodized SWG-assisted CDC.
Conclusion